OpenStudy (anonymous):

-48 ÷ (7 – 9)3 – 2[1 – 5(2 – 6) + 32]. simplify and show work

OpenStudy (anonymous):

its not 32 its 3^2

OpenStudy (anonymous):

-48÷(-2)3-2[1-5(-4)+9] =-48÷(-6)-2[1+20+9] =8-2[30] =8-60=-52